The units of a certain number is one greater than twice the tens digit. If the digits are reversed, the new number is 36 more than the original number. Find the original number
Answer by jorel1380(3719)

Let xy be the number.
2x+1=y
10x+y+36=10y+x
10x+2x+1+36=10(2x+1)+x
12x+37=21x+10
9x=27
x=3
y=7
The original number is 37..